OpenStudy (anonymous):

I was proving d/dx (e^x). Our teacher mentioned the proof, but I have a question. Anyone ?

OpenStudy (anonymous):

\[\LARGE \lim_{h \rightarrow 0}~\frac{f(x+h)-f(x)}{h}\]\[\LARGE \lim_{h \rightarrow 0}~\frac{e^{x+h}-e^{x}}{h}\]\[\LARGE \lim_{h \rightarrow 0}~\frac{e^{x}e^{h}-e^{x}}{h}\]\[\LARGE \lim_{h \rightarrow 0}~\frac{e^{x}(e^{h}-1)}{h}\]\[\LARGE e^{x} \times \lim_{h \rightarrow 0}~\frac{(e^{h}-1)}{h}\]

OpenStudy (anonymous):

Why is the last limit, \[\Large \color{red}{\lim_{h \rightarrow 0}~\frac{(e^{h}-1)}{h}}\] equal to 1 ?

OpenStudy (amistre64):

what is e^0 - 1?

OpenStudy (anonymous):

yes, so that is just saying that 0/0 = 1 ?

OpenStudy (anonymous):

I mean e^0-1 = 1-1 =0 So the limit is 0/0, and that is 1. That's what you want to say ?

OpenStudy (ikram002p):

use LH rule

OpenStudy (anonymous):

Haha, I haven't proved e^x yet

OpenStudy (anonymous):

I won't be able to know e^h like this, this will become an endless cycle, see ?

OpenStudy (amistre64):

thats what i was leaning towards yes, 0/0 limits to 1

OpenStudy (anonymous):

I was always thinking that it 0/0 is undefined. But isn't there anything better to this than saying that 0/0=1 ?

OpenStudy (ikram002p):

no :O

OpenStudy (anonymous):

I looked up on yahoo, but all I can see is L'H's. Well, I am familiar with it, despite that I'm only taking Calc1, but even using L'H's I am unable to get the limit, because although when I derive the bottom, the "h" (w/ respect to h) I get 1, I will have to derive e^h on the top and that is not proven yet (for me) to do this.

OpenStudy (ikram002p):

well the limit is 0/0 thus ur allowed to use LH rule when limit f(x)/g(x) as x goes to 0 is 0/0 then limit f(x)/g(x) as x goes to 0=limit f'(x)/g'(x) as x goes to 0

OpenStudy (anonymous):

No, L'H's doesn't work for my given (above) reason.

OpenStudy (anonymous):

Math sucks !

OpenStudy (anonymous):

(apparently)

OpenStudy (amistre64):

e^h = 1 + h + h^2/2! + h^3/3! + .... e^h - 1 = h + h^2/2! + h^3/3! + .... (e^h-1)/h = 1 + h/2! + h^2/3! + ... when h=0, 1+ 0 + 0 + 0 + ....

OpenStudy (anonymous):

ooh, cool. I kind of heard that, but I never thought that there will be a need for that. It so nicely intervenes over here! Tnx armistre!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Sorry, can't write a second testimonial -:(

OpenStudy (accessdenied):

The series expansion also uses the derivative of e^x for the coefficients, right? Or can it come through other means that I just didn't think of? :P

OpenStudy (amistre64):

well, that is one way to develop the series yes. but if its already known before hand

OpenStudy (anonymous):

well, the coefficient if you mean something like d/dx e^(2x) lim f(x+h) - f(x) h approaches 0 ------------ h lim e^(2(x+h)) - e^(2x) h approaches 0 ------------ h lim e^(2x)e^(2h) - e^(2x) h approaches 0 ------------------ h lim e^(2x) (1 - e^(2h) ) h approaches 0 ------------------ h lim (1 - e^(2h) ) e^(2x) times h approaches 0 ------------- h e^(2h) - 1^2 = h^2 +.... e^(2h) - 1 = h^2 +.... when h=0, 1+ 0 + 0 + 0 + ....

OpenStudy (accessdenied):

Sorry, I wanted to post earlier but OS bonked out and I had to go. Got it, although I guess I just don't find it convincing enough personally! One thing that I was thinking about was implicit differentiation by taking the natural log. of both sides. But I figured maybe you wanted the limit definition proof specifically, or the natural log derivative was also off-limits. Here was something I stumbled across that looked promising: https://www.physicsforums.com/threads/solving-for-derivative-of-e-x-using-limit-definition.161965/#post-1280952 It is a bit more involved but the actions seem sound. Basically just using the substitutions of variables and definition of e to make the simplifications necessary.

OpenStudy (accessdenied):

And no, by coefficients I meant the coefficients of the expanded series form of e^x. \( \displaystyle e^x = \sum_{n=0}^{\infty} \color{green}{\dfrac{1}{n!} } x^n\) You can derive the coefficients by repeated differentiation. e^x = a_0 + a_1 x + a_2 x^2 + ... x=0 e^0 = a_0 + 0 + 0 + ... = a_0 = 1 d/dx(e^x) = d/dx(a_0 + a_1 x + a_2 x^2 + ... ) e^x = a_1 + 2a_2 x + .3a_3 x^2 + ... x=0 e^0 = a_1 + 0 + 0 + 0... = a_1 = 1. Due to power rule, you can see that the next will be 2a_2 = 1, or 1/2 = 1/2!. Then 3*2 a_3 = 1, or 1/3/2 = 1/3!. And so on. The only problem is that you're using derivatives of e^x again, unless you already have this form given to you.

OpenStudy (xapproachesinfinity):

the proof of derivatives of lnx and e^x involve some given limits related to definition of e and uses the ln inverse of e^x to obtain the result without series expansion

OpenStudy (xapproachesinfinity):

you may want to check MIT videos to see how they did it i think the prof that is lecturing proved it

OpenStudy (xapproachesinfinity):

you can prove it using implicit differentiation y=e^x x=lny given we know d/dx(lnx) you can prove it easily using the definition